OpenStudy (rayman5766):

What is the value of p – (–3.5) when p = 8.2?

OpenStudy (rayman5766):

also I have 10 more

OpenStudy (texaschic101):

p - (-3.5) when p = 8.2 now we sub 8.2 - (-3.5) --- negative times a negative is a positive, thus changing your problem to 8.2 + 3.5 = 11.7

OpenStudy (rayman5766):

oh ok hold

OpenStudy (rayman5766):

its -6.85

OpenStudy (rayman5766):

What value of y makes the equation true? 3.4 + y = –6.85 A. –10.25 B. –3.25 C. 3.25 D. 10.25